OpenStudy (anonymous):

find limit lim 4 / x>5 x^2 -25

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5}\frac{ 4 }{ x ^{2}-25 }=\lim_{x \rightarrow 5}\frac{ 0 }{2x }=\frac{ 0 }{ 2\times5 }=\frac{ 0 }{10 }=0\] (here i have differentiated the numerator & denominator.

OpenStudy (anonymous):

Thank you!